OpenStudy (anonymous):

Factor: 3x^2 + 11x + 6

OpenStudy (ganpat):

so have u did factorization before ?

OpenStudy (anonymous):

Hint: 18 has factors as 9 and 2.. You can change the middle one according to it..

OpenStudy (anonymous):

@waterineyes ok

OpenStudy (anonymous):

Can you do now @jasminecrump18

OpenStudy (ganpat):

3x^2 + 11x + 6 3x2 + 9x +2x + 6 = 0 as, 9 * 2 = 18 & 9 + 2 = 11 3x (x +3) + 2 (x+ 3) = 0 (3x + 2) (x+3) = 0 x = -2/3 and x = -3...

OpenStudy (anonymous):

thnks @Ganpat i understand it now

OpenStudy (ganpat):

glad, and please do reply if someone asks u question.. Y / N .. :) :)

OpenStudy (anonymous):

ok